OpenStudy (anonymous):

Evaluate cot 45° without using a calculator by using ratios in a reference triangle.

OpenStudy (wolfe8):

It's by drawing a right-angled triangle with sides 1,1 and 2 with a line cutting it in 2 I believe(I forgot). You can look up reference triangle on Google.

OpenStudy (jdoe0001):

without using a calculator, but using your Unit Circle :)

OpenStudy (jdoe0001):

\(\bf cot(\theta) = \cfrac{cos(\theta)}{sin(\theta)}\)

OpenStudy (jdoe0001):

notice in your Unit Circle that \(\bf cos(45^o) = \cfrac{\sqrt{2}}{2} \qquad \qquad sin(45^o) = \cfrac{\sqrt{2}}{2}\\\quad \\ cot(45^o) = \cfrac{\quad \frac{\sqrt{2}}{2}\quad }{\frac{\sqrt{2}}{2}} \implies \square?\)

OpenStudy (anonymous):

thank u guys

OpenStudy (jdoe0001):

yw